OpenStudy (anonymous):

Solve for x: 2(cube root of)(10-3x)=(cube root of)(2-x) Please show the steps! :)

OpenStudy (anonymous):

2(cube root of)(10-3x)=(cube root of)(2-x) 2(10-3x)^(1/3)=(2-x)^1/3 put both sides to the power of 3, we get (2^3)(10-3x)=(2-x) 8(10-3x)=(2-x)

OpenStudy (anonymous):

80-24x=2-x 78=23x x=78/23

OpenStudy (anonymous):

did you get the idea? ant question?..lol..

OpenStudy (anonymous):

Nope! That's the same answer as I had too! I wasn't too sure on the 2^3 thing or leave it 2. Thank you again! :)